A woman is six times as old her daughter, Three years ago, she was 1 times as old as her daughter,find their ages
Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

daughter's age=x. Note, 3 years earlier, she had to be more than 6 times older, so 1 doesn't work. I will use 11.
woman's=6x
6x-3=11(x-3)
6x-3=11x-33
-5x=-30
x=6 years old now, mother is 36.
3 years ago, daughter 3 years old, mother 33, 11 times older.

Answer by MathTherapy(10552)


A woman is six times as old her daughter, Three years ago, she was 1 times as old as her daughter,find their ages
IMPOSSIBLE! Try AGAIN